The length of a rectangle is twice the width. The perimeter is 24 meters. Find the length and the width! Please and thank you!
Thanks a bunch!!!

Answer by piglet162431(90)

formula= 24=2l+2w
If l=2w, then 2w(2)=4w
24=4w+2w
24=6w
4=w
Since 4 is the width, length is 8.
4+4+8+8= 8+16=24